# GNN-DT: A Graph Neural Network Enhanced Decision Transformer for Efficient Optimization in Dynamic Environments

### v2

@@ A.3 The MDP State Space
Table 6 shows a comparison of key EV charging metrics for the 25-station problem after 100 evaluations, including heuristic algorithms, Charge As Fast as Possible (CAFAP) and BaU, and DT variants with the optimal solution, which assumes future knowledge. GNN-DT shows remarkable performance, achieving a close approximation to the optimal solution, particularly in user satisfaction (99.3% ± 0.03%) and power violation (21.7 ± 22.8 kW). It outperforms both BaU and DT variants in terms of energy discharged, power violation, and costs. Notably, GNN-DT performs well even compared to Q-DT, while maintaining competitive execution time, albeit slightly slower than the simpler models. The results underscore the effectiveness of GNN-DT in managing complex EV charging tasks, demonstrating its potential for real-world applications where future knowledge is not available.
@@
{tblr}
cells = c,t, vline2 = 2-70.05em, hline1,8 = -0.08em, hline2 = -0.05em, Algorithm & Energy
@@
CAFAP $1.3$$0.2$ $0.00$ ±$0.00$ $100.0$$0.0$ $1289.2$$261.8$ $-277$$165$ $-1.974$$0.283$ $0.001$
BaU $1.3$$0.2$ $0.00$ ±$0.00$ $99.9$$0.2$ $10.5$$9.4$ $-255$$156$ $-0.679$$0.067$ $0.001$
DT $0.9$$0.1$ $0.03$$0.01$ $94.4$$1.6$ $58.7$$28.3$ $-173$$104$ $-0.462$$0.093$ $0.006$
Q-DT $1.0$$0.1$ $0.00$ ±$0.00$ $93.6$$2.1$ $20.1$$21.4$ $-187$$113$ $-0.665$$0.135$ $0.010$
GNN-DT (Ours) $0.9$$0.1$ $0.19$$0.03$ $99.3$$0.2$ $21.7$$22.8$ $-142$$89$ $-0.027$$0.023$ $0.023$
Optimal (Offline) $1.9$$0.2$ $1.08$$0.19$ $99.1$$0.2$ $2.0$$4.6$ $-119$$84$ $-0.020$$0.015$
Figure 6: Complete results for a single evaluation scenario:(a) illustrating individual EV battery trajectories across all 25 charging stations, and (b) presenting the actual aggregate power usage against the power limit for the entire EV fleet.
Fig. 6 presents the results from a single evaluation scenario, focusing on the performance of various charging strategies across all 25 charging stations. Fig. 6a illustrates the individual battery trajectories for each EV across the stations, showing how the actual SoC evolves over time. The desired SoC is compared against the results from different algorithms: BaU, GNN-DT, DT, Q-DT, and the Optimal (Offline) solution. It is evident that GNN-DT closely tracks the desired SoC across all stations, outperforming the other methods, particularly in terms of maintaining the target SoC. Fig. 6b provides insights into the aggregate power usage of the entire EV fleet, where the actual power used is compared to the setpoint power. GNN-DT closely aligns with the power setpoint, demonstrating effective power management, while other methods such as BaU and Q-DT show greater deviations, indicating less efficient power usage. These results underline the superior performance of GNN-DT in optimizing charging strategies while adhering to power constraints.
A.5 State Transition Distribution Experiments
To analyze the generalization capability of our model, we examine how it performs under environments with varying state transition probabilities. Specifically, we visualize key environment variables that directly impact the transition dynamics. Figure 7a–d presents the probability distributions of EV arrival time, departure time, duration of stay, and state of SoC at arrival across four scenarios: the original training environment and environments with small, medium, and extreme variations. These plots help quantify the extent of variation in each case. Additionally, Figure 7e illustrates the temporal distribution of the power limit in each scenario, providing further insight into the differences in environment configuration.
(a) Arrival time
(b) Departure time
(c) Time of stay
(d) SoC at arrival
(e) Power limit
@@ A.5 State Transition Distribution Experiments
Figure 7: Overview of the five key metrics: (a) arrival time, (b) departure time, (c) power limit, (d) state of charge upon arrival, and (e) time of stay.
Appendix B Appendix: Training
Here, we present the hyperparameter settings used for training DT, Q-DT, and GNN-DT, accompanied by detailed training curves that illustrate the convergence of each model.
B.1 Training Hyperparameters
Table 7: Algorithm hyperparameters.
Hyperparameter Value
Batch Size $64$ (for large scale), $128$ (for small scale)
Learning Rate $10^{-4}$
Weight Decay $10^{-4}$
Number of Steps per Iteration $1000$
Number of Decoder Layers $3$
Number of Attention Heads $4$
Embedding Dimension $256$
GNN Embedder Feature Dimension $16$
GNN Hidden Dimension $32$ (for small scale), $64$ (for large scale)
Number of GCN Layers $3$
Number of Epochs $250$ (for small scale), $400$ (for large scale)
Number of Steps per Iteration $1000$ (for small scale), $3000$ (for large scale)
Embedding Dimension $128$ (for small scale), $256$ (for large scale)
CPU Memory (GB) $8$ (for small scale), $16$ or $40$ (for large scale)
Time Limit (hours) $10$ (for small scale), $20$ or $46$ (for large scale)
B.2 Impact of larger context lengths (K)
Figure 8: GNN-DT performance for larger context lengths (K).
B.3 Detailed Training Curves
Fig. 9 provides a detailed comparison of training curves for various algorithm-dataset-context length ($K$) combinations, highlighting the significant impact of the training sample size on performance. The figure includes training results for DT, Q-DT, and GNN-DT across different datasets (Optimal, Random, and BaU) and context lengths ($K=2$ and $K=10$), with each plot showing the performance for 100, 1,000, and 10,000 trajectory samples.
Figure 9: Complete comparison of training curves for combinations of algorithms-K-training dataset for different numbers of samples.
B.4 Mixed Dataset Training Curves
Fig. 10 presents the learning curves for the mixed datasets approach, comparing performance across various combinations of Optimal, Random, and BaU datasets for both $K=2$ and $K=10$. Fig. 10a illustrates the results for mixed Optimal datasets, where different proportions of the Optimal and Random datasets (e.g., 50% Optimal + 50% Random) are used for training. The performance becomes more unstable as more Random data is included. Interestingly, the performance for all the mixed datasets demonstrates better maximum reward reached compared to the Optimal-only dataset, highlighting the benefits of combining high-quality and lower-quality data. Fig. 10b and Table 8 shows similar trends for the Mixed-BaU datasets. While the BaU dataset alone performs worse than the Optimal dataset, mixing it with Random data still yields improvements, with the 75% BaU and 25% Random combination showing the best results. The results underscore the potential of mixing datasets to improve training performance, especially when high-quality data (Optimal or BaU) is supplemented with lower-quality Random data.
(a) Mixed-Optimal
(b) Mixed-BaU
Figure 10: Mixed datasets learning curves for optimal and BaU.
Dataset Total Traj. Avg. Dataset Reward GNN-DT Reward (×$\times 10^{5}$)
K=2 k=10
Random (Rnd.) 100% $1000$ $-2.37$$0.39$ $-0.863$ $-1.187$
BaU 25% + Rnd. 75% $1000$ $-1.93$$0.80$ $-0.578$ $-0.461$
BaU 50% + Rnd. 50% $1000$ $-1.51$$0.87$ $-0.665$ $\mathbf{-0.447}$
BaU 75% + Rnd. 25% $1000$ $-1.09$$0.76$ $\mathbf{-0.421}$ $-0.471$
BaU 100% $1000$ $-0.01$$0.01$ $-0.654$ $-0.572$

### v3

@@ 6.5 Component Ablation Study
GNN-DT shows remarkable performance, achieving a close approximation to the optimal solution, particularly in user satisfaction (99.3% ± 0.03%) and power violation (21.7 ± 22.8 kW). It outperforms both BaU and DT variants in terms of energy discharged, power violation, and costs. Notably, GNN-DT performs well even compared to Q-DT, while maintaining competitive execution time, albeit slightly slower than the simpler models. The results underscore the effectiveness of GNN-DT in managing complex EV charging tasks, demonstrating its potential for real-world applications where future knowledge is not available.
@@
Algorithm Energy
@@
CAFAP $1.3$ ±$0.2$ $0.00$ ±$0.00$ $100.0$ ±$0.0$ $1289.2$ ±$261.8$ 277-277 ±$165$ 1.974-1.974 ±$0.283$ $0.001$
BaU $1.3$ ±$0.2$ $0.00$ ±$0.00$ $99.9$ ±$0.2$ $10.5$ ±$9.4$ 255-255 ±$156$ 0.679-0.679 ±$0.067$ $0.001$
DT $0.9$ ±$0.1$ $0.03$ ±$0.01$ $94.4$ ±$1.6$ $58.7$ ±$28.3$ 173-173 ±$104$ 0.462-0.462 ±$0.093$ $0.006$
Q-DT $1.0$ ±$0.1$ $0.00$ ±$0.00$ $93.6$ ±$2.1$ $20.1$ ±$21.4$ 187-187 ±$113$ 0.665-0.665 ±$0.135$ $0.010$
GNN-DT (Ours) $0.9$ ±$0.1$ $0.19$ ±$0.03$ $99.3$ ±$0.2$ $21.7$ ±$22.8$ 142-142 ±$89$ 0.027-0.027 ±$0.023$ $0.023$
Optimal (Offline) $1.9$ ±$0.2$ $1.08$ ±$0.19$ $99.1$ ±$0.2$ $2.0$ ±$4.6$ 119-119 ±$84$ 0.020-0.020 ±$0.015$ -
6.7 Illustrative Example of EV Charging
With the models trained, we proceed to compare the behavior of the best baseline models trained (DT, Q-DT, GNN-DT) against the heuristic BaU algorithm in an EV charging scenario. Fig. 3(a) presents the SoC progress for three EVs connected one after the other to a single charger throughout the simulation, while Fig. 3(b) shows the actions of all chargers taken by each algorithm. At the beginning of the simulation, EVs arrive at the charging station with unknown initial SoCs. Upon connection, they communicate their departure times and desired SoC levels to the CPO. Leveraging this information, along with real-time electricity price signals and power constraints, each algorithm determines optimal charging and discharging actions.
In Fig. 3(a), the heuristic BaU algorithm consistently overcharges the EVs, often exceeding the desired SoC levels. In contrast, both DT and Q-DT fail to satisfy the desired SoC. Conversely, GNN-DT successfully achieves the desired SoC for all EVs, closely mirroring the behavior of the optimal algorithm. This demonstrates GNN-DT’s ability to precisely control charging based on dynamic state information. Fig. 3(b) provides further insights into the actions taken by each algorithm. The optimal solution primarily employs maximum charging or discharging power, since it knows the future. In comparison, GNN-DT exhibits a more refined approach, modulating charging power within a range of -6 to 11 kW. Baseline DT and Q-DT display a narrower range of actions, limiting their ability to optimize the charging schedules and adapt to varying conditions. These results underscore the superior capability of GNN-DT in managing the complexities of EV charging dynamics.
6.8 Generalization and Scalability Analysis
@@
Figure 5: Overview of the five key state transition variables across different scenarios: (a) arrival time, (b) departure time, (c) power limit, (d) state of charge upon arrival, and (e) time of stay.
Evaluating the generalization of RL models across varying state transition probabilities is crucial for ensuring consistent performance under diverse conditions Wang et al. (2020). To evaluate the generalization capabilities of GNN-DT, three additional environments with different state transition probabilities are designed. The key environment variables that directly impact the state transition dynamics are visualized in Figure 6.1. In detail, Figure 6.1a–d presents the probability distributions of EV arrival time, departure time, duration of stay, and state of SoC at arrival across four scenarios: the original training environment and environments with small, medium, and extreme variations. These plots help quantify the extent of variation in each case. Additionally, Figure 6.1e illustrates the temporal distribution of the power limit in each scenario, providing further insight into the differences in environment configuration.
In Fig. 5(a), the generalization capabilities of GNN-DT and other baselines are assessed in environments with small, medium, and extreme variations in state transition probabilities. While the baseline methods experience significant performance drops as the evaluation environment deviates from the training setting, GNN-DT maintains strong performance across all scenarios. This highlights the critical role of GNN-based embeddings in improving model robustness and generalization. A key advantage of the GNN-DT architecture, not present in classic DTs, is its invariance to problem size, i.e., the same RL agent can be applied to both smaller and larger-scale environments. Fig. 5(b) illustrates the scalability and generalization performance of GNN-DT compared to the BaU algorithm and Optimal policy. GNN-DT, trained on a 25-charger setup, was tested on 5, 50, 75, and 100-charger environments. While its performance predictably declines at larger scales, since it wasn’t trained on them, it still outperforms the BaU heuristic, demonstrating robustness to problem-size variation. Training GNN-DT on a mix of charger numbers could potentially further improve its adaptability.
The scalability and effectiveness of GNN-DT were tested when trained on a significantly larger optimization problem involving 250 charging stations. In this scenario, the model must handle up to 250 action variables per step and over 1,000 state variables, which include critical information such as power limits and battery levels. The results presented in Table 8 demonstrate that GNN-DT shows promise for addressing more complex optimization tasks. However, the model requires a substantial increase in both the number of training trajectories and memory resources to maintain efficiency, highlighting a well-known limitation of DT-based approaches. Scaling the problem 10× roughly multiplies GPU memory usage, e.g. storing 3,000 trajectories takes $\approx 2$ GB for 25 chargers versus $\approx 20$ GB for 250. While this can bottleneck large‐scale training, parallelization and mini‐batching mitigate it, and overall compute scales with the transformer’s context length K (see Fig. 3), pointing to interesting directions for very large problem graphs as future work.
Total
Trajectories
Avg. Dataset
Reward
GNN-DT
Reward
Random $3000$ 22.39-22.39 ±$1.49$ 9.34-9.34
BaU $3000$ 6.67-6.67 ±$0.32$ 4.23-4.23
Optimal $3000$ 0.08-0.08 ±$0.03$ $\mathbf{-0.27}$
7 Conclusions
Acknowledgements
The study was partially funded by the DriVe2X research and innovation project from the European Commission with grant numbers 101056934. The authors acknowledge the use of computational resources of the DelftBlue supercomputer, provided by Delft High Performance Computing Centre (https://www.tudelft.nl/dhpc). This work used the Dutch national e-infrastructure with the support of the SURF Cooperative using grant no. EINF-5716.
[…]